# Resolving the Ti–V Phase Diagram Discrepancy with First-Principles Calculations and Bayesian Learning

### v1

@@ Abstract
Conflicting experiments disagree on whether the titanium–vanadium (Ti–V) binary alloy exhibits a body-centred cubic (BCC) miscibility gap or remains completely soluble. A leading hypothesis attributes the miscibility gap to oxygen contamination during alloy preparation. To resolve this controversy, we use an ab initio + machine-learning workflow that couples an actively-trained Moment Tensor Potential to Bayesian thermodynamic inference. Using this workflow, we obtain Ti–V binary system across the entire composition range, together with confidence intervals in the thermodynamic limit $(N\rightarrow\infty)$. The resulting diagram reproduces all experimental features, demonstrating the robustness of our approach, and clearly favors the variant with a BCC miscibility gap terminating at $T_{\mathrm{crit}}=980\;\text{K}$ and $c_{\mathrm{crit}}=0.67$. Because oxygen was excluded from simulations, the gap cannot be attributed to impurity effects, contradicting recent CALPHAD reassessments.
I Introduction
Phase diagrams are foundational tools in materials science, serving as guides for designing and synthesizing materials under specific thermodynamic conditions. In this work, we study the phase diagram of the titanium–vanadium (Ti–V) binary system, which has been a subject of long-standing conflicting interpretations in the literature: one report a body-centred cubic (BCC) miscibility gap (Figure 1, version A), and the other with complete solid solubility across the entire composition range (Figure 1, version B).
The experimental study by Nakano [1] indicates formation of a miscibility gap for temperatures above 948 K. However, the experimental investigations by Khaled [2, 3] and Fuming [4] indicate a complete solid solubility. The most recent CALPHAD (CALculation of PHAse Diagrams) reassessment of phase diagram by Lindwall (2018) [5] also indicate complete solid solubility and attribute the miscibility gap formation to the presence of oxygen impurities.
Modern machine-learning interatomic potentials (MLIPs), especially systematically improvable families such as moment tensor potentials (MTPs) [6], now deliver ab initio accuracy at a fraction of the computational cost. Leveraging these advances, we constructed a MLIP-accelerated ab initio Ti-V phase diagram to resolve the conflicting experimental interpretations. Because the simulation conditions are fully controlled, potential sources of contamination—such as oxygen impurities—can be systematically eliminated.
@@ I Introduction
Using a Bayesian learning framework, we infer the Ti–V phase diagram across the full temperature–composition range and report confidence intervals for every phase boundary. The resulting diagram reproduces all experimental features, demonstrating the robustness of our approach, and clearly favors the variant with a BCC miscibility gap, showing that oxygen impurities are unnecessary for gap formation.
@@
In this study, we construct the Ti–V phase diagram through the following workflow. (i) An active-learning protocol based on the D-optimality criterion selects ab initio atomic configurations, from which we fit an interatomic potential. (ii) Using this potential, we generate a thermodynamic data set—with associated computational uncertainties—for every equilibrium phase: the body-centred cubic (BCC) solid, the hexagonal-close-packed (HCP) solid, and the liquid. (iii) From these data we reconstruct free-energy surfaces and locate the corresponding phase-transition boundaries. The workflow is readily transferable to other alloy systems — especially those where experimental information is sparse or conflicting.
@@
The remainder of the paper is organized as follows. Section II reviews related work. Section III presents the Bayesian algorithm for phase-diagram inference. Section IV.1 details the construction of the interatomic potential. Finally, Section IV.2 describes the generation of the thermodynamic data set and the inference of the Ti–V phase diagram.
II Related Work
@@
Two first-principles investigations have explored the Ti–V phase diagram, each adopting distinct approximations. Sluiter et al.[7] combined a tight-binding Hamiltonian — whose hopping integrals are fitted once to linear augmented-plane-wave (LAPW) band structures — with the Coherent Potential Approximation (CPA) and the Generalized Perturbation Method (GPM). More recently, Chinnappan et al. [8] employed a cluster expansion parametrized on density functional theory calculations, combined with Monte Carlo sampling. To capture vibrational effects, they used the bond stiffness versus bond length method, which assumes a linear relationship between bond length and the nearest-neighbour force constant.
@@
III Methods
@@
The remainder of this section is organized as follows. Section III.1 introduces the Gaussian process regression framework. The subsequent subsections describe how different data sources are incorporated into the algorithm: free-energy asymptotics (Section III.2), molecular dynamics data (Section III.3), phonon calculations (Section III.4), and melting points (Section III.5). Section III.6 focuses specifically on the Bayesian estimation of the melting temperature and its associated confidence interval. Finally, Section III.7 presents the functional form of the moment-tensor potential employed in this study.
@@
III.1 Gaussian Process Regression
@@
III.2 Free energy asymptotics
@@
III.3 Molecular Dynamics Data
@@ III.4 Phonon Calculations
For the liquid phase the additive constant is determined from the melting temperature obtained via $NPT$ solid–liquid coexistence simulations, as described in Sec. III.5.
III.5 Melting Points
Thanks to the Bayesian approach, our algorithm incorporates the data on melting points along with their corresponding uncertainties. Melting temperatures are evaluated by the coexistence algorithm summarized in Sec. III.6, following the procedure of Ref. [15].
@@ III.5 Melting Points
III.6 Bayesian Coexistence Algorithm for Melting Point Prediction
@@
III.7 Moment Tensor Potential
@@
Here, $T^{(\beta)}(|\mathbf{r}_{ij}|)$ denotes the Chebyshev polynomial of degree $\beta$, defined on the interval $[R_{\mathrm{min}},R_{\mathrm{cut}}]$, and the term $(R_{\mathrm{cut}}-|\mathbf{r}_{ij}|)^{2}$ ensures a smooth decay to zero at the cutoff boundary.
@@
IV Results
IV.1 Development of the Moment Tensor Potential
[…]
@@ IV.1 Development of the Moment Tensor Potential
IV.2 Phase Diagram Workflow
IV.2.1 Exploratory sampling and surrogate model
@@
Using the supplementary dataset, we trained a gradient boosting regressor (Scikit-learn implementation [21]) to capture the dependence of the concentration $c$ on $\mu$, $T$, and $N$. The resulting surrogate model enabled an almost uniform sampling of the $T$$c$$N$ space, yielding 62 data points in BCC region, 28 in HCP, and 36 in the liquid phase (Figure 4); these constitute the points for the MD simulations described below.
IV.2.2 Thermodynamic Dataset
@@
The MD dataset provides training data for the derivatives of the entropy with respect to temperature and concentration. Knowing these derivatives, we reconstructed the absolute entropy by anchoring each phase at a reference point. To obtain this reference, we carried out phonon calculations, evaluating the quantity $\frac{1}{N}\log\det\hat{H}$, which gives the entropy of the HCP and BCC phases at $T=0K$. We then employed Gaussian-process regression to extrapolate $\frac{1}{N}\log\det\hat{H}$ to the thermodynamic limit $N\to\infty$, as illustrated in Figure 5(a). The extrapolated values are $\frac{1}{N}\log\det\hat{H}=4.23165\pm 5\cdot 10^{-5}$ for HCP titanium, $\frac{1}{N}\log\det\hat{H}=5.94765\pm 6\times 10^{-5}$ for BCC vanadium.
[…]
@@
IV.2.3 Free‐energy reconstruction
@@ IV.2.4 Ti–V phase diagram
Our algorithm produced the phase diagram of the Ti–V binary system across the entire composition range, together with $2\sigma$ confidence intervals in the thermodynamic limit $(N\rightarrow\infty)$, as shown in Figure 8. Experimental data available in the literature [24, 25, 26, 27, 28, 23] are plotted for comparison. The algorithm reproduces the key experimental features: the Ti-rich boundary of the HCP–BCC coexistence region is matched quantitatively, whereas the V-rich boundary and the solidus–liquidus envelope are captured qualitatively. Moreover, the predicted HCP–BCC transition temperature, $T_{\mathrm{HCP\!-\!BCC}}=1156\pm 10K$, matches the experimental value of $1155K$ within one standard deviation.
@@
The Ti–V phase diagram predicted in this work exhibits a BCC miscibility gap in the V-rich region, for which no experimental data are currently available. The critical point is located at $T_{\mathrm{crit}}=980\pm 3K$ and $c_{\mathrm{crit}}=0.673\pm 0.005$. Notably, the phase diagram was generated entirely from quantum-mechanical calculations, with no experimental input. This enables an unbiased comparison with the two experimental assessments reported in the literature (Figure 9). Our results favor the version A, which includes a BCC miscibility gap, over Version B, which do not. Moreover, the present findings do not support the claim of Lindwall et al. [5] that oxygen impurities are responsible for the formation of the miscibility gap: the composition in our simulations was strictly controlled, and oxygen was not included.
V Summary and Concluding Remarks
Two experimental assessments of the Ti–V phase diagram predominate in the literature: Version A reports a miscibility gap (e.g., Murray 1989 [11]), whereas Version B shows no such gap (e.g., Lindwall 2018 [5]). Here, we constructed the Ti–V phase diagram over the full temperature–composition range entirely from quantum-mechanical calculations. A molecular-dynamics dataset for the BCC, HCP, and liquid phases—supplemented by zero-kelvin phonon calculations and melting points of Ti and V obtained from coexistence simulations—served as input to our Bayesian learning framework. The algorithm propagates the uncertainty in the input data to the phase-transition boundaries and extrapolates them to the thermodynamic limit ($N\!\to\!\infty$), yielding an uncertainty-quantified Ti–V phase diagram. When compared with the available experimental data, our algorithm quantitatively reproduced the Ti-rich boundary of the HCP–BCC two-phase region and qualitatively captures both the V-rich boundary and the solidus–liquidus envelope. The predicted HCP–BCC transition temperature, $T_{\mathrm{HCP\!-\!BCC}}=1156\pm 10K$, matches the experimental value of $1155K$ within one standard deviation.
@@ V Summary and Concluding Remarks
This study highlights the power of our ab initio, uncertainty-aware workflow for constructing phase diagrams, which is capable of resolving discrepancies in experimental phase-diagram assessments. In the future, we plan to extend our Bayesian framework to three-component alloys, which will, for instance, allow a direct quantification of the influence of oxygen on Ti–V phase stability and an explicit treatment of oxygen-bearing phases.
V.1 Acknowledgments

### v2

# Clarifying the Ti–V Phase Diagram Using First-Principles Calculations and Bayesian Learning

@@ Abstract
Conflicting experiments disagree on whether the titanium–vanadium (Ti–V) binary alloy exhibits a body-centred cubic (BCC) miscibility gap or remains completely soluble. A leading hypothesis attributes the miscibility gap to oxygen contamination during alloy preparation. To resolve this disagreement, we use an ab initio + machine-learning workflow that couples an actively-trained Moment Tensor Potential with Bayesian inference of free energy surface. This workflow enables construction of the Ti–V phase diagram across the full composition range with systematically reduced statistical and finite-size errors. The resulting diagram reproduces all experimental features, demonstrating the robustness of our approach, and clearly favors the variant with a BCC miscibility gap terminating at $T_{\mathrm{crit}}=980\;\text{K}$ and $c_{\mathrm{crit}}=0.67$. Because our simulations model a perfectly oxygen-free Ti–V system, the observed gap cannot originate from impurity effects, in contrast to recent CALPHAD reassessments.
keywords:
phase diagram , free energy , Bayesian learning
journal: Computational Materials Science
[…]
{graphicalabstract}[Uncaptioned image]
{highlights}
Ti–V phase diagram constructed with ab initio accuracy using actively trained machine-learning interatomic potential.
Statistical and finite-size errors of phase transition boundaries are systematically eliminated.
Computed Ti–V phase diagram reproduces all experimental features and indicates a body-centered cubic miscibility gap.
Findings challenge the earlier hypothesis attributing the miscibility gap to oxygen impurities.
1 Introduction
Phase diagrams are foundational tools in materials science, serving as guides for designing and synthesizing materials under specific thermodynamic conditions. Recent advances in machine-learned interatomic potentials [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] now enable molecular simulations with ab initio accuracy at a fraction of the computational cost. Despite these significant advances, constructing phase diagrams from first principles remains challenging.
The most accurate method for phase diagram construction is thermodynamic integration [11, 12, 13], which reconstructs the free energy from molecular dynamics (MD) simulations by integrating its derivatives with respect to simulation parameters along a path to a reference state. While being very accurate, this method requires multiple MD simulations to discretize each integration path, making it computationally expensive to map the entire free energy surface. Another widely used approach is coexistence simulations [14, 15, 16], where phases are simulated in direct contact. However, because such simulations must be repeated for each set of conditions, they become impractical for systems with many phase transition boundaries.
A recently proposed Bayesian framework [17, 18] retains the strengths of free-energy reconstruction while avoiding the drawbacks of both approaches. MD simulations are performed only once for each selected temperature and composition, while systematically eliminating finite-size effects and propagating statistical uncertainties from the MD data into the predicted free energies and phase boundaries. This approach produces the full phase diagram while substantially reducing the number of required simulations.
In this work, we apply the Bayesian framework to the titanium–vanadium (Ti–V) binary system, a case with long-standing experimental controversy. One set of studies [19, 20] identifies a body-centred cubic (BCC) miscibility gap (Figure 1, version A), while others [21, 22, 23] claim complete solid solubility across the entire composition range (Figure 1, version B). The most recent CALPHAD (CALculation of PHAse Diagrams)[24] reassessment by Lindwall [25] also indicates complete solid solubility, attributing earlier reports of a miscibility gap to oxygen impurities.
@@ 1 Introduction
We infer the complete temperature–composition Ti-V phase diagram with confidence intervals for each phase boundary using the Bayesian framework. The resulting diagram reproduces all experimental features, demonstrating the robustness of our approach, and clearly favors the variant with a BCC miscibility gap, showing that oxygen impurities, absent by construction in our computations, are unnecessary for gap formation.
@@
Our workflow proceeds in three main stages: (i) an active-learning protocol based on the D-optimality criterion selects ab initio atomic configurations, from which we fit an interatomic potential; (ii) we employ the fitted potential to generate a thermodynamic data set — with associated computational uncertainties — for each equilibrium phase: the body-centred cubic (BCC) solid, the hexagonal-close-packed (HCP) solid, and the liquid; (iii) from these data, we reconstruct free-energy surfaces and locate the corresponding phase-transition boundaries. The workflow is readily transferable to other alloy systems — especially those where experimental information is sparse or conflicting.
@@
The remainder of the paper is organized as follows. Section 2 reviews related work. Section 3 presents the Bayesian algorithm for phase-diagram inference. Section 4.1 details the construction of the interatomic potential. Finally, Section 4.2 describes the generation of the thermodynamic data set and the inference of the Ti–V phase diagram.
2 Related Work
@@
Two first-principles investigations have explored the Ti–V phase diagram, each adopting distinct approximations. Sluiter et al.[26] combined a tight-binding Hamiltonian — whose hopping integrals are fitted once to linear augmented-plane-wave (LAPW) band structures — with the Coherent Potential Approximation (CPA) and the Generalized Perturbation Method (GPM), using the Cluster Variational Method (CVM) to compute configurational free energies. More recently, Chinnappan et al. [27] employed a cluster expansion parametrized on density functional theory calculations, combined with Monte Carlo sampling. To capture vibrational effects, they used the bond stiffness versus bond length method, which assumes a linear relationship between bond length and the nearest-neighbour force constant.
@@
We also note the recent experimental studies of Smith et al. [30] and MacLeod et al. [31] on commercial titanium alloys (e.g., Ti-6Al-4V, Ti-36Nb-2Ta-0.3O) that investigate the effects of oxygen impurities on phase transition boundaries and thermodynamic properties of these alloys.
3 Methods
@@
The remainder of this section is organized as follows. Section 3.1 introduces the Gaussian process regression framework. The subsequent subsections describe how different data sources are incorporated into the algorithm: free-energy asymptotics (Section 3.2), molecular dynamics data (Section 3.3), phonon calculations (Section 3.4), and melting points (Section 3.5). Section 3.6 focuses specifically on the Bayesian estimation of the melting temperature and its associated confidence interval. Finally, Section 3.7 presents the functional form of the moment-tensor potential employed in this study.
@@
3.1 Gaussian Process Regression
@@
3.2 Free energy asymptotics
@@
3.3 Molecular Dynamics Data
@@ 3.4 Phonon Calculations
For the liquid phase the additive constant is determined from the melting temperature obtained via $NPT$ solid–liquid coexistence simulations, as described in Sec. 3.5.
3.5 Melting Points
Thanks to the Bayesian approach, our algorithm incorporates the data on melting points along with their corresponding uncertainties. Melting temperatures are evaluated by the coexistence algorithm summarized in Sec. 3.6, following the procedure of Ref. [34].
@@ 3.5 Melting Points
3.6 Bayesian Coexistence Algorithm for Melting Point Prediction
@@
3.7 Moment Tensor Potential
@@
Here, $T^{(\beta)}(|\mathbf{r}_{ij}|)$ denotes the Chebyshev polynomial of degree $\beta$, defined on the interval $R_{\mathrm{min}},R_{\mathrm{cut}}]$, and the term $(R_{\mathrm{cut}}-|\mathbf{r}_{ij}|)^{2}$ ensures a smooth decay to zero at the cutoff boundary.
@@
4 Results
4.1 Development of the Moment Tensor Potential
We trained the machine-learning potential on configurations obtained from density-functional-theory (DFT) calculations performed with the Vienna Ab initio Simulation Package (VASP) [35] using the projector-augmented-wave (PAW) method [36]. The Perdew–Burke–Ernzerhof generalized-gradient approximation (PBE-GGA) [37] was employed as the exchange–correlation functional. The self-consistent cycle was converged to an electronic-energy tolerance of $3\times 10^{-5}$ eV with a Gaussian smearing width of 0.2 eV. The ENCUT parameter was chosen as 466 eV which is $1.7\cdot\max(\texttt{ENMAX\_TI},\texttt{ENMAX\_V})$, where ENMAX is the energy cutoff of the corresponding pseudopotential, and a $\Gamma$-centered k-point mesh having a spacing of 0.16 Å-1 was used to ensure convergence of DFT computations up to 1 meV/at.
@@ 4.1 Development of the Moment Tensor Potential
4.2 Phase Diagram Workflow
4.2.1 Exploratory sampling and surrogate model
@@
Using the supplementary dataset, we trained a gradient boosting regressor (Scikit-learn implementation [40]) to capture the dependence of the concentration $c$ on $\mu$, $T$, and $N$. The resulting surrogate model enabled an almost uniform sampling of the TTccNN space, yielding 62 data points in BCC region, 28 in HCP, and 36 in the liquid phase (Figure 5); these constitute the points for the MD simulations described below.
4.2.2 Thermodynamic Dataset
@@
The MD dataset provides training data for the derivatives of the entropy with respect to temperature and concentration. Knowing these derivatives, we reconstructed the absolute entropy by anchoring each phase at a reference point. To obtain this reference, we carried out phonon calculations, evaluating the quantity $\frac{1}{N}\log\det\hat{H}$, which gives the entropy of the HCP and BCC phases at $T=0K$. We then employed Gaussian-process regression to extrapolate $\frac{1}{N}\log\det\hat{H}$ to the thermodynamic limit $N\to\infty$, as illustrated in Figure 4(a).
We still require an anchor point for the liquid phase, which we obtain from the melting temperatures. At the melting point, the free energies of the solid and liquid phases are equal; because the solid free energies are already known, this equality furnishes the liquid anchor. To determine the melting temperatures, we employ the algorithm of Klimanova et al. [34], which combines nonlinear Bayesian regression with finite-size extrapolation to the thermodynamic limit and provides a confidence interval for the resulting melting points. The workflow is illustrated in Figure 4(b) for BCC titanium: first, melting temperatures and their uncertainties are estimated for several system sizes, and then Gaussian-process regression is used to extrapolate to $N\!\to\!\infty$. We obtain $T_{\mathrm{m}}(\mathrm{Ti})=1763\pm 2K$ and $T_{\mathrm{m}}(\mathrm{V})=2180\pm 2K$. The vanadium value agrees well with the experimental melting point of $2187~\text{K}$ [42], whereas the titanium value is underestimated by roughly $200~\text{K}$—a common bias of DFT calculations with the PBE functional [43, 44] (experimental $T_{\mathrm{m}}(\mathrm{Ti})=1941~\text{K}$ [42]).
@@
4.2.3 Free‐energy reconstruction
@@ 4.2.4 Ti–V phase diagram
Our algorithm produced the phase diagram of the Ti–V binary system across the entire composition range, together with $2\sigma$ confidence intervals in the thermodynamic limit $(N\rightarrow\infty)$, as shown in Figure 8. Experimental data available in the literature [45, 46, 47, 48, 49, 42] are plotted for comparison. (comparation with previous first principles computations The algorithm reproduces the key experimental features: the Ti-rich boundary of the HCP–BCC coexistence region is matched quantitatively, whereas the V-rich boundary and the solidus–liquidus envelope are captured qualitatively. Moreover, the predicted HCP–BCC transition temperature, $T_{\mathrm{HCP\!-\!BCC}}=1156\pm 10K$, matches the experimental value of $1155K$ within one standard deviation.
@@
The Ti–V phase diagram predicted in this work exhibits a BCC miscibility gap in the V-rich region, for which no experimental data are currently available. The critical point is located at $T_{\mathrm{crit}}=980\pm 3K$ and $c_{\mathrm{crit}}=0.673\pm 0.005$. Notably, the phase diagram was generated entirely from quantum-mechanical calculations, with no experimental input for comparison with previous first-principles studies [26, 27], see A). This enables an unbiased comparison with the two experimental assessments reported in the literature (Figure 9). Our results favor the version A, which includes a BCC miscibility gap, over Version B, which do not. Moreover, the present findings do not support the claim of Lindwall et al. [25] that oxygen impurities are responsible for the formation of the miscibility gap: the composition in our simulations was strictly controlled, and oxygen was not included.
5 Summary and Concluding Remarks
Two experimental assessments of the Ti–V phase diagram predominate in the literature: Version A reports a miscibility gap (e.g., Murray 1989 [20]), whereas Version B shows no such gap (e.g., Lindwall 2018 [25]); both are shown in Figure 9. Here, we constructed the Ti–V phase diagram over the full temperature–composition range entirely from quantum-mechanical calculations. A molecular-dynamics dataset for the BCC, HCP, and liquid phases—supplemented by zero-kelvin phonon calculations and melting points of Ti and V obtained from coexistence simulations—served as input to our Bayesian learning framework. The algorithm propagates the uncertainty in the input data to the phase-transition boundaries and extrapolates them to the thermodynamic limit ($N\!\to\!\infty$), yielding an uncertainty-quantified Ti–V phase diagram. When compared with the available experimental data, our algorithm quantitatively reproduced the Ti-rich boundary of the HCP–BCC two-phase region and qualitatively captures both the V-rich boundary and the solidus–liquidus envelope. The predicted HCP–BCC transition temperature, $T_{\mathrm{HCP\!-\!BCC}}=1156\pm 10K$, matches the experimental value of $1155K$ within one standard deviation.
@@ 5 Summary and Concluding Remarks
Beyond resolving this specific discrepancy, our workflow highlights particular compositions and temperatures where experiments would be most informative. For example, our predictions place Ti-70V at 800 K within the miscibility gap region, indicating that targeted experiments on high-purity samples at this composition would provide a direct test of the contested feature of the Ti–V system.
Looking ahead, we aim to extend our ab initio, uncertainty-aware Bayesian framework to ternary alloys, enabling, in particular, direct quantification of oxygen’s influence on Ti–V phase stability and explicit treatment of oxygen-bearing phases.
Acknowledgments
@@
Data Availability
The data that support the findings of this study are openly available in “Ti–V Moment Tensor Potential Dataset” at https://doi.org/10.5281/zenodo.15755826, reference number [50]. The repository provides both the reference dataset and the fitted MTP for the titanium–vanadium binary alloy. All reference energies and forces were generated with density functional theory calculations, while the potential itself was trained via active learning.
Appendix A Comparison of First-Principles Ti-V Phase Diagrams
Figure 10: Computed phase diagram of Ti-V compared to two interpretations of experimental data (Version A and Version B). The plot also presents two theoretical studies, both utilizing ab initio computations to parametrize the interactions in the system. Sluiter [26] utilizes the cluster variation method to compute the free energy of the solid phase, while Chinnappan [27] uses Monte Carlo for compositional sampling (along with the bond stiffness versus bond length method to capture the vibrational contribution).